For u = (5,-4) and v= (-1, 2) find u•v.
Answer #1

Collar u = (5,-4) V=(-1,2) U.de = (5,-4). (-1,2) =(5)(-1) +64)(2) --5-O (a,.b. ). 19, b) a q + b by uon=-13 And